# LaViPlan : Language-Guided Visual Path Planning with RLVR

### v2

@@ Abstract
Out-of-distribution (OOD) scenarios in autonomous driving refer to situations that deviate from the training domain, often leading to unexpected and potentially hazardous behavior from planners that lack prior exposure to such cases. Recently, Vision-Language Models (VLMs) have been introduced into autonomous driving research for their promising generalization capabilities in OOD settings. Early studies demonstrated that VLMs could recognize OOD scenarios and generate user-level decisions such as “go straight” or “turn right.” However, a new challenge has emerged due to the misalignment between the VLM’s high-level decisions or visual reasoning expressed in language, and the low-level predicted trajectories interpreted as actions. In this paper, we propose LaViPlan, a framework that leverages Reinforcement Learning with Verifiable Rewards (RLVR) to optimize VLMs using planning-oriented metrics. This approach addresses the vision-language-action misalignment observed in existing VLMs fine-tuned via supervised learning, which can recognize driving scenarios but often produce context-unaware decisions. Experimental results demonstrate that our method improves situational awareness and decision-making under OOD conditions, highlighting its potential to mitigate the misalignment issue. This work introduces a promising post-training paradigm for VLM agents in the context of autonomous driving.
@@ 1 Introduction
Out-of-distribution (OOD) scenarios in autonomous driving refer to rare or novel situations that deviate from the training domain, often leading to unexpected and unsafe behavior by learned planning policies. These scenarios pose critical challenges, particularly when the planner fails to generalize beyond its supervised experience. To address this, recent research has explored the integration of Vision-Language Models (VLMs) [4, 15, 1] into autonomous driving. VLMs have demonstrated strong generalization capabilities across diverse tasks and modalities, making them a promising approach for handling OOD scenarios. Early research showed that VLMs could identify anomalous or unseen driving contexts and generate high-level user-aligned decisions (e.g., “go straight,” “turn right”) [10, 11]. However, while VLMs can recognize and describe OOD scenes, their final decisions—especially in the form of predicted trajectories—can often be misaligned with the visual reasoning they produce. This issue reflects a broader challenge in aligning language-based reasoning with action-level planning, which we refer to as the vision-language-action misalignment. In this paper, we propose to address the misalignment by leveraging Reinforcement Learning with Verifiable Rewards (RLVR), guided by planning-oriented metrics. Our method aims to steer VLMs toward context-aware decision-making that is consistent with their situational reasoning. Our key contributions are summarized as follows:
  • We propose a reinforcement learning framework that explicitly optimizes planning-oriented metrics in VLMs. Our method demonstrates a step toward aligning language-guided reasoning with action-oriented tasks in autonomous driving.
@@ 2 Related Works
2.1 Vision and Language Foundation Models
End-to-end autonomous driving has recently gained attention due to its simplicity and efficiency, as well as its ability to address the suboptimality issues that arise from misaligned objectives between modular components. This paradigm reduces information loss and computational overhead by eliminating intermediate representations. However, the visual abstraction in end-to-end systems often simplifies complex scene information, which can result in the loss of critical cues. Furthermore, achieving generalization across diverse driving scenarios remains challenging. Simply increasing the scale of training data is not always feasible, particularly in long-tail or out-of-distribution (OOD) scenarios where labeled data is limited or sparse. To overcome these challenges, recent studies have explored integrating Vision-Language Models (VLMs) into autonomous driving, as VLMs are capable of grounding and reasoning about previously unseen situations through their multi-modal understanding. These models typically combine pre-trained large language models (LLMs) with visual encoders, where driving instructions or ego vehicle states are provided as textual input to the LLM, and single- or multi-view images are passed through the visual encoder [12, 10, 19, 25, 22, 30, 31, 23]. More recent research has extended these models with techniques such as 3D perceptual positional embeddings [27, 6, 34] and counterfactual learning [27, 20] to enable context-aware decision-making in complex environments.
2.2 Reinforcement Learning for Language Models
Reinforcement Learning from Human Feedback (RLHF) has demonstrated the effectiveness of aligning VLMs with human preferences with Proximal Policy Optimization (PPO) [13]. However, RLHF has computational overhead due to the multiple components, including a reference model, reward model, critic, and a newly trained generative model. Additionally, the reliance on human annotators further increases the cost of training. In contrast, Group Relative Policy Optimization (GRPO) [8] enhances the reasoning and arithmetic capabilities of VLMs by learning directly from comparisons between the policy model and the reference model, without requiring an explicit reward model or critic. The authors advocate for a rule-based reward model, arguing that neural reward models may be vulnerable to reward hacking during large-scale reinforcement learning. Reinforcement fine-tuning with RLVR has recently been applied to tasks such as object detection, classification [16], mathematics [5], and code generation, where ground-truth outputs are either unambiguous or can be verified through test cases. It has also been extended to robotics domains, including manipulator control [24, 28] and autonomous driving [18, 11], where the correctness of the output can be evaluated through physical simulations or pre-defined behavioral criteria. The objective nature of these tasks makes the design and validation of reward signals relatively straightforward. The effectiveness of GRPO has also been theoretically analyzed: recent work shows that GRPO improves the success rate of reward-maximizing outputs through a fixed-point dynamic process and provides guarantees on preference amplification over the reference model [17]. Furthermore, its alignment mechanism differs fundamentally from RLHF’s logarithmic pooling, instead applying nonlinear transformations that prioritize relatively more rewarding outputs under group-based comparisons [26]. In this work, we explore the use of Multimodal Large Language Models (MLLMs) for autonomous driving by leveraging RLVR with planning-oriented metrics, such as Average Displacement Error (ADE) and Final Displacement Error (FDE). Our approach aims to fine-tune MLLMs through objective and verifiable reward signals, demonstrating the effectiveness of RLVR in planning-centric tasks such as trajectory prediction in autonomous driving.
@@ 3.1 Preliminary
Reinforcement Learning from Verifier Rewards RLVR aims to optimize a policy $\pi_{\theta}$ by maximizing the expected reward from a verifier while maintaining proximity to a reference policy $\pi_{\text{ref}}$ to prevent over-optimization. The objective function is formulated as:
@@
where $q$ represents the input query, $o$ denotes the generated output, $R(q,o)$ is the reward function provided by the verifier, and $\beta$ is a regularization coefficient that controls the trade-off between reward maximization and KL divergence penalty. The KL divergence term ensures that the learned policy does not deviate significantly from the reference policy, maintaining stability during training.
Group Relative Policy Optimization GRPO extends the Proximal Policy Optimization (PPO)[21] by incorporating group-relative advantage estimation. The advantage function is computed using group statistics to reduce variance:
@@
where $G$ represents the group size, $R_{i}$ is the reward for the $i$-th response, and the advantage $\hat{A}_{i,t}$ is normalized using the mean and standard deviation of rewards within the group. This normalization helps stabilize training by reducing the impact of reward scale variations.
@@ 3.2 RLVR with Planning-Oriented Metrics
In this work, we apply GRPO to autonomous driving by designing a reward function that directly reflects planning performance as shown in Fig. 2. Specifically, we define the reward using commonly adopted trajectory evaluation metrics: Average Displacement Error (ADE) and Final Displacement Error (FDE).
@@ 4 Experiment
In-Domain Dataset. The ROADWork dataset [7] serves as an OOD benchmark focused on road construction scenarios. Among the entire dataset, 5,430 samples contain image, scene description, and a list of trajectories in image plane. To demonstrate that reinforcement fine-tuning can yield performance improvements even with data less than supervised fine-tuning, we divided the dataset into two subsets: 4,344 samples for supervised fine-tuning and 1,086 samples (20% of full samples) for reinforcement fine-tuning. Training details for supervised fine-tuning and reinforcement fine-tuning can be seen in the Appendix A.
Subsets Defined by Trajectory Variance. We split the dataset into two subsets based on the variance of x-coordinates in the trajectory. One subset comprises trajectories with low lateral variance representing straight trajectory (Easy). In contrast, the other subset comprises trajectories with higher lateral variance and more curved trajectories involving left and right turns (Hard). Details of the dataset splitting procedure are provided in the Appendix B.
Out-Domain Dataset. The CODA-LM [2] extends the CODA dataset [14] which includes various corner cases such as road construction and adverse weather conditions by augmenting it with natural language captions. The generalization ability of our method under OOD conditions was qualitatively assessed using the corner-case dataset not limited to construction zones.
Dataset and Prompt Design for Reinforcement Fine-Tuning. The dataset for reinforcement fine-tuning consists of a reasoning description enclosed within <think></think> tags, which includes visual reasoning process from the image, and a predicted trajectory represented as a list of 20 (x, y) image coordinates enclosed within <answer></answer> tags as shown in Tab. 1. Specifically, the <think></think> section provides the rationale behind the prediction, while the <answer></answer> section contains the autonomous vehicle’s future trajectory as a sequence of coordinate pairs.
@@
4.1 Experimental Results
Table 2: Comparison of model performance on ROADWork dataset. ADE and FDE are presented as percentages by normalizing with image resolution for better readability. The models with supervised fine-tuning were trained on the full dataset of 5K samples, while LaViPlan was fine-tuned with 4K samples for supervised learning and an additional 1K samples for reinforcement fine-tuning. Bolded values indicate the best performance.
@@ 4.1.1 Evaluation Results
4.1.2 From Linguistic Consistency to Functional Reasoning.
@@
SFT RFT Change
BERTScore 0.748 0.712 -4.81%
Entailment (%) 44.8 37.7 -15.85%
Neutral (%) 25.3 29.1 +15.02%
Contradiction (%) 29.9 33.1 +10.70%
Figure 3: Qualitative comparison of scene reasoning between the SFT (top) and RFT (bottom) models across six scenarios in [7]. While SFT tends to produce more verbose, human-like descriptions that closely resemble natural annotations, the RFT model emphasizes core, task-relevant elements such as cones, barriers, and work vehicles. This shift in reasoning reflects a transition from linguistic fidelity to planning-oriented abstraction, enabling the model to better prioritize hazards and navigable space—even at the cost of similarity to ground-truth descriptions.
We evaluate the linguistic effects of our method on the combined easy and hard subsets using BERTScore [33] and Natural Language Inference (NLI) [3], as shown in Tab. 3. Compared to supervised fine-tuning (SFT), reinforcement fine-tuning (RFT) leads to a decrease in BERTScore, as well as a reduction in entailment and an increase in neutral and contradiction labels. While these results indicate a divergence from ground-truth human-written descriptions, they do not necessarily imply degraded reasoning ability. Rather, we observe a meaningful shift in the model’s reasoning style. As shown in Fig. 3, scene descriptions from RFT tend to be more concise and hazard-focused emphasizing critical elements such as barriers, cones, and work vehicles in terms of planning. In contrast, SFT tends to replicate verbose and linguistically rich but often functionally redundant information. In several scenarios, RFT omits less relevant descriptors (e.g., sidewalks or drum lines) in favor of spatial cues that better support trajectory-level decisions. This trade-off suggests that conventional language metrics may underrepresent reasoning utility in autonomous driving contexts. Despite lower alignment with human-like phrasing, RFT yields improved planning performance, indicating that more effective and actionable reasoning can emerge even as linguistic similarity declines. We advocate for a shift toward task-aware evaluation of vision-language models in safety-critical domains, where functional relevance should take precedence over semantic mimicry.
4.1.3 Evaluating Out-of-Distribution Generalization.
We qualitatively evaluate the proposed method on CODA-LM [2], a zero-shot dataset composed of diverse corner cases involving out-of-distribution (OOD) road hazards. Since CODA-LM does not provide ground-truth trajectories, we rely on visual inspection of the predicted paths to assess plausibility, safety, and contextual awareness. As shown in Fig. 4, the model fine-tuned with RLVR produces more realistic and contextually appropriate trajectories compared to the one trained via supervised fine-tuning. In scenarios involving roadwork, blocked lanes, and ambiguous path constraints, the supervised model often generates overly linear or risk-prone trajectories that lack appropriate deviation or caution. In contrast, the proposed method consistently generates smoother, safer, and more context-aware trajectories, often adjusting path curvature to avoid obstacles such as cones, barriers, and vehicles. These results demonstrate the model’s capacity to generalize to previously unseen situations by aligning its reasoning and path generation with latent planning cues in the scene, even without explicit supervision. While the absence of quantitative metrics in CODA-LM limits numerical evaluation, these qualitative outcomes suggest that reinforcement fine-tuning enhances the model’s ability to adapt to novel driving contexts.

### v4

@@ Abstract
Out-of-distribution (OOD) scenarios in autonomous driving pose critical challenges, as planners often fail to generalize beyond their training experience, leading to unsafe or unexpected behavior. Vision-Language Models (VLMs) have shown promise in handling such scenarios by providing high-level scene understanding and user-aligned decisions. However, existing VLMs often exhibit a misalignment between their language-based reasoning and the low-level trajectories required for action-level planning. In this paper, we propose LaViPlan, a framework that leverages Reinforcement Learning with Verifiable Rewards (RLVR) to fine-tune VLMs using planning-oriented metrics. Experimental results show that LaViPlan improves planning performance across both in-domain and out-of-domain datasets. While linguistic fidelity slightly decreases after RLVR-based fine-tuning, qualitative evaluation indicates that the outputs remain coherent. We also conduct ablation studies to analyze the effects of sampling ratio and reasoning guidance, highlighting how these design choices influence performance. These findings demonstrate the potential of RLVR as a post-training paradigm for aligning language-guided reasoning with action-level planning in autonomous driving.
@@ 1 Introduction
Out-of-distribution (OOD) scenarios in autonomous driving refer to rare or novel situations that deviate from the training domain, often leading to unexpected and unsafe behavior by learned planning policies. These scenarios pose critical challenges, particularly when the planner fails to generalize beyond its supervised experience.
To address this, recent research has explored the integration of VLMs [5, 19, 1] into autonomous driving. VLMs have demonstrated strong generalization capabilities across diverse tasks and modalities, making them a promising approach for handling OOD scenarios. Early research showed that VLMs could identify unseen driving contexts and generate high-level decisions [14, 15].
However, while VLMs can recognize and describe OOD scenes, their final decisions—especially in the form of predicted trajectories—can often be misaligned with the visual reasoning they produce. This issue reflects a broader challenge in aligning language-based reasoning with action-level planning, which we refer to as the vision-language-action misalignment.
To mitigate misalignment issues in other domains, GRPO-based reinforcement fine-tuning has been proposed, showing particular effectiveness in improving performance on numerical prediction tasks [20, 30]. By leveraging multiple candidate outputs and preference-based rewards instead of relying solely on a single ground-truth target, GRPO provides richer supervisory signals that enhance both accuracy and generalization beyond the training distribution. This success suggests that GRPO could be a promising approach to address the misalignment in VLM for autonomous driving.
Building on this insight, we propose to address the misalignment by leveraging Reinforcement Learning with Verifiable Rewards (RLVR), where planning-oriented metrics serve as verifiable reward signals. Our method aims to steer VLMs toward context-aware decision-making that is consistent with their situational reasoning. Our key contributions are summarized as follows:
  • We propose a reinforcement learning framework that explicitly optimizes planning-oriented metrics in VLMs, demonstrating a step toward aligning language-guided reasoning with action-oriented planning tasks in autonomous driving.
@@ 2 Related Works
2.1 VLMs for Autonomous Driving
End-to-end autonomous driving has recently attracted attention due to its simplicity, efficiency, and ability to mitigate suboptimality arising from misaligned objectives between modular components. By eliminating intermediate representations, this paradigm reduces information loss and computational overhead. However, visual abstraction in end-to-end systems can oversimplify complex scene information, potentially discarding critical cues. Achieving robust generalization across diverse driving scenarios also remains challenging, especially in long-tail scenarios where labeled data is limited or sparse.
To address these limitations, recent studies have explored integrating Vision-Language Models (VLMs) into autonomous driving. VLMs leverage multi-modal understanding to reason about previously unseen situations. Typically, these models combine pre-trained large language models (LLMs) with visual encoders: driving instructions or ego vehicle states are fed as textual input to the LLM, while single- or multi-view images are processed by the visual encoder [16, 14, 25, 31, 28, 35, 36, 29]. Recent extensions incorporate 3D perceptual positional embeddings [33, 7, 39] and counterfactual learning [33, 26] to enable context-aware decision-making in complex driving environments.
2.2 Preference Learning for Alignment
While VLMs demonstrate strong multi-modal reasoning capabilities, their outputs may still be misaligned with downstream action-level tasks, such as trajectory prediction. One common approach to address this issue is Reinforcement Learning from Human Feedback (RLHF), which aligns VLMs with human preferences using Proximal Policy Optimization (PPO) [27, 17, 12]. However, RLHF requires multiple components—including a reference model, reward model, critic, and a newly trained generative model—and relies heavily on human-in-the-loop feedback, resulting in substantial computational and labor costs. This limitation is particularly critical for precise tasks like path planning in autonomous driving.
To overcome these challenges, Group Relative Policy Optimization (GRPO) [10] has been proposed. GRPO enhances VLMs’ reasoning and arithmetic capabilities by learning directly from comparisons between the policy model and a reference model, without requiring an explicit reward model or critic. The authors advocate for rule-based reward signals, arguing that neural reward models may be vulnerable to reward hacking during large-scale reinforcement learning. Unlike standard or preference-based RL [27, 23], GRPO directly optimizes verifiable, task-specific metrics. This aligns the policy with downstream objectives, such as path planning, rather than linguistic mimicry, and reduces both human supervision and computational cost.
Building on GRPO, Reinforcement Learning with Verifiable Reward (RLVR) has been successfully applied to tasks where correctness can be objectively evaluated, including object detection, classification [20], mathematics [6], and code generation. RLVR has also been extended to robotics, such as manipulator control [30, 34] and autonomous driving [22, 15], where reward signals can be verified via simulations or pre-defined behavioral criteria. Theoretical analyses show that RLVR with GRPO improves the success rate of reward-maximizing outputs through a dynamic process converging to a fixed point and guarantees preference amplification over the reference model [21, 32].
In this work, we extend RLVR with GRPO to autonomous driving by leveraging planning-oriented metrics, such as Average Displacement Error (ADE) and Final Displacement Error (FDE), as verifiable reward signals. Our approach fine-tunes VLMs to produce trajectories that are consistent with situational reasoning, improving both out-of-distribution generalization and alignment between language-based planning and action-level execution.
@@ 3.1 Preliminary
Reinforcement Learning with Verifiable Rewards. RLVR aims to optimize a policy $\pi_{\theta}$ by maximizing the expected reward while maintaining proximity to a reference policy $\pi_{\text{ref}}$ to prevent over-optimization. The objective function is formulated as:
@@
where $q$ represents the input query, $o$ denotes the generated trajectory, $R(q,o)$ is the reward function, and $\beta$ is a regularization coefficient that controls the trade-off between reward maximization and KL divergence penalty. The KL divergence term ensures that the learned policy does not deviate significantly from the reference policy, maintaining stability during training.
Group Relative Policy Optimization. GRPO extends the Proximal Policy Optimization (PPO)[27] by incorporating group-relative advantage estimation. The advantage function is computed using group statistics to reduce variance:
@@
where $G$ represents the number of generated trajectories, i.e. the group size, $R_{i}$ is the reward for the $i$-th trajectory, and the advantage $\hat{A}_{i,t}$ is normalized using the mean and standard deviation of rewards within the group. This normalization helps stabilize training by reducing the impact of reward scale variations.
@@ 3.2 RLVR with Planning-Oriented Metrics
In this work, we apply GRPO to autonomous driving by designing a reward function that directly reflects planning performance, as shown in Fig. 2. We build upon the insight that GRPO enables learning pairwise preferences among multiple candidate trajectories. Unlike single-reward updates, this approach exposes the policy to richer and more diverse supervisory signals, providing multiple perspectives on the range of candidate trajectories. This diversity encourages the policy to learn robust behaviors that are not overfitted to specific situations, which ultimately improves generalization across a wide range of scenarios. Therefore, we define the reward using commonly adopted trajectory evaluation metrics: ADE and FDE.
@@ 4 Experiment
In-Domain Dataset. The ROADWork dataset [9] serves as an OOD benchmark focused on road construction scenarios. Among the entire dataset, 5,430 samples contain image, scene description, and a list of trajectories in image plane. To demonstrate that reinforcement fine-tuning can yield performance improvements even with data less than supervised fine-tuning, we divided the dataset into two subsets: 4,344 samples for supervised fine-tuning and 1,086 samples (20% of full samples) for reinforcement fine-tuning. Details for training can be seen in the Appendix A.
Furthermore, inspired by KITTI [8], we split the dataset into two subsets based on the variance of x-coordinates in the trajectory. One subset comprises trajectories with low lateral variance representing straight trajectory (Easy). In contrast, the other subset comprises trajectories with higher lateral variance and more curved trajectories involving left and right turns (Hard). Details of the dataset splitting procedure are provided in the Appendix B.
Out-of-Domain Dataset. The CODA-LM [2] extends the CODA dataset [18] which includes various corner cases such as road construction and adverse weather conditions by augmenting it with natural language captions. The generalization ability of our method under OOD conditions was assessed using the corner-case dataset not limited to construction zones.
What is the best baseline model? We choose Qwen2VL-2B-Instruct as our baseline. This decision is supported by existing studies [20, 15, 3] that integrate RLVR with diverse tasks and consistently adopt this model, thereby demonstrating its validity and robustness in multimodal learning contexts.
@@
Dataset and Prompt for Reinforcement Fine-Tuning. The dataset for reinforcement fine-tuning consists of a reasoning description enclosed within <think></think> tags, which includes visual reasoning process from the image, and a predicted trajectory represented as a list of image coordinates enclosed within <answer></answer> tags as shown in Tab. 1. Specifically, the <think></think> section provides the rationale behind the prediction, while the <answer></answer> section contains the future trajectory as a sequence of coordinate pairs.
4.1 In-domain Evaluation
Table 2: Comparison of model performance on ROADWork dataset. ADE and FDE are presented as percentages by normalizing with image resolution for better readability. The models with supervised fine-tuning were trained on the full dataset of 5K samples, while LaViPlan was fine-tuned with 4K samples for supervised learning and an additional 1K samples for reinforcement fine-tuning. Bolded values indicate the best performance and N/A means no available result.
@@ 4.1 In-domain Evaluation
Domain-Specific Models
Senna N/A N/A N/A N/A
DriveLM (w/ LLaMA-Adapter) 37.10 38.40 56.99 56.90
@@
Domain-Specific Models
Senna 5.71 5.73 6.58 7.46
DriveLM (w/ LLaMA-Adapter) 6.73 7.79 6.87 8.43
@@ 4.1.1 Overall Results
4.1.2 From Linguistic Consistency to Functional Reasoning
Figure 3: Qualitative comparison of scene reasoning between the SFT (top) and LaViPlan (bottom) models across six scenarios in [9]. While SFT tends to produce more verbose, human-like descriptions that closely resemble natural annotations, LaViPlan emphasizes core, task-relevant elements such as cones, barriers, and work vehicles. This shift in reasoning reflects a transition from linguistic fidelity to planning-oriented abstraction, enabling the model to better prioritize hazards and navigable space—even at the cost of similarity to ground-truth descriptions.
We evaluate the linguistic effects of our method on the combined easy and hard subsets using BERTScore [38] and Natural Language Inference (NLI) [4], as shown in Tab. 3. Compared to supervised fine-tuning (SFT), reinforcement fine-tuning (LaViPlan) leads to a decrease in BERTScore, as well as a reduction in entailment and an increase in neutral and contradiction labels. While these results indicate a divergence from ground-truth human-written descriptions, they do not necessarily imply degraded reasoning ability. Rather, we observe a meaningful shift in the model’s reasoning style. As shown in Fig. 3, scene descriptions from LaViPlan tend to be more concise and hazard-focused emphasizing critical elements such as barriers, cones, and work vehicles in terms of planning.
In contrast, SFT tends to replicate verbose and linguistically rich but often functionally redundant information. In several scenarios, LaViPlan omits less relevant descriptors (e.g., sidewalks or drum lines) in favor of spatial cues that better support trajectory-level decisions. This trade-off suggests that conventional language metrics may underrepresent reasoning utility in autonomous driving contexts. Despite lower alignment with human-like phrasing, LaViPlan yields improved planning performance, indicating that more effective and actionable reasoning can emerge even as linguistic similarity declines. We advocate for a shift toward task-aware evaluation of vision-language models in safety-critical domains, where functional relevance should take precedence over semantic mimicry.
@@ 4.1.2 From Linguistic Consistency to Functional Reasoning
SFT LaViPlan $\Delta$
BERTScore $\uparrow$ 0.748 0.712 -4.81%
Entailment (%) $\uparrow$ 44.8 37.7 -15.85%
Neutral (%) $\downarrow$ 25.3 29.1 +15.02%
Contradiction (%) $\downarrow$ 29.9 33.1 +10.70%
4.2 Out-of-domain Evaluation
We evaluate the proposed method on CODA-LM [2], a zero-shot dataset composed of diverse corner cases involving out-of-distribution (OOD) road hazards. Since CODA-LM does not provide ground-truth trajectories, we evaluate the method based on 2D bounding box in the dataset.
4.2.1 Evaluation Metrics
To quantitatively assess the safety of the predicted trajectories, we evaluate three core indicators that capture a distinct aspect of safety. The detailed procedure for obtaining F, C, and P can be found in the Appendix C:
  • Fail Rate (F): A trajectory is considered a failure if it invades any 2D bounding box in the scene. This reflects how often unsafe behaviors occur.
  • Collision Count (C): The average number of bounding boxes that are violated per trajectory across all scenes. This captures the extent of interaction with multiple objects
  • Penetration Length (P): When a collision occurs, this metric measures the extent of trajectory penetration into the bounding box area, providing a proxy for the severity of the failure. This quantifies the depth of intrusion, emphasizing the potential danger in collision.
Inspired by [13], we compute an overall safety score by applying min-max normalization to each metric, followed by a weighted average:
$\text{SafetyScore}_{i}=\sum_{j\in\{F,C,P\}}w_{j}\cdot\left(1-\frac{x_{ij}-\min_{k}x_{kj}}{\max_{k}x_{kj}-\min_{k}x_{kj}}\right)$ (7)
The safety score for each model $i$ is computed as a weighted sum over metrics $j\in\{F,C,P\}$ — failure rate, collision count, and penetration length — where each metric is min-max normalized across all models $k$. Then, we employ multiple weighting schemes to capture different evaluation perspectives. This approach enables a balanced and comprehensive analysis of model robustness and trade-offs under various out-of-distribution scenarios.
  • Balanced ($w_{f}$ = 0.4, $w_{c}$ = 0.3, $w_{p}$ = 0.3): This evaluation reflects a balanced trade-off between safety and driving success under general OOD scenarios.
  • Safety-Focused ($w_{f}$ = 0.3, $w_{c}$ = 0.2, $w_{p}$ = 0.5): Each indicator is first min-max normalized to the [0, 1] range, and then aggregated using predefined weights. This approach highlights the relative performance differences across models and is useful for identifying overall safety trends.
  • Performance-Focused ($w_{f}$ = 0.5, $w_{c}$ = 0.3, $w_{p}$ = 0.2): This scheme prioritizes driving success by assigning a high weight to the trajectory failure rate, while giving smaller weights to collision count and penetration length, respectively. It emphasizes task completion while still accounting for safety.
  • Equal Weight ($w_{f}$ = 0.33, $w_{c}$ = 0.33, $w_{p}$ = 0.34): All three safety indicators are weighted equally (approximately 33% each). This evaluation provides a neutral assessment without emphasizing any specific aspect, making it suitable for overall robustness analysis.
Table 4: Safety scores under different weighting schemes for failure (F), collision (C), and penetration (P).
[…]
Baseline 0.40 0.30 0.50 0.33
SFT (5k) 0.60 0.59 0.56 0.63
LaViPlan 0.64 0.73 0.56 0.70
Table 4 presents the safety scores of each model under four different weighting schemes. Across all evaluation settings, the proposed method consistently achieves the highest scores, indicating superior safety performance in both driving success and physical impact mitigation. In particular, it shows a significant advantage under the Safety-Focused, minimizing penetration depth and collisions. Although SFT (fine-tuned with whole dataset) performs relatively well, especially under the Performance-Focused, it lags behind LaViPlan in overall safety. The baseline model consistently scores the lowest, confirming its limited capability in handling OOD scenarios safely. The qualitative results of SFT and ours can be found in Appendix D.